Given market demand Qd = 50 - P, and market supply P = Qs + 5




A) Find the market equilibrium price and quantity?




B) What would be the state of the market if market price was fixed at Birr 25 per unit?




C) Calculate and interpret price elasticity of demand at the equilibrium point.

1
Expert's answer
2021-11-08T16:44:16-0500

A) Find the market equilibrium price and quantity?

At equilibrium "Qd=Qs"

Therefore, "P=Qs+5=Qd+5=50-P+5"

"P+P=55"

"P=\frac{55}{2}=27.5"

"Q=50-P=50-27.5=22.5"

B) What would be the state of the market if market price was fixed at Birr 25 per unit?

 If "P=25, Qd=50-25=25"

"Qs=P-5=25-5=20"

So, quantity demanded is higher than quantity supplied.

Thus, there is excess demand in the market.

C) Calculate and interpret price elasticity of demand at the equilibrium point.

 The price elasticity of demand= "\frac{dQd}{dP}\times \frac{P}{Qd}"

"Qd=50-P"

ie, "\frac{dQd}{dP}=-1"

Price elasticity of demand at equilibrium"=-1\times\frac{27.5}{22.5}=-1.22"